OpenStudy (anonymous):

Solve for x: -2(x + 3) = -2(x + 1) - 4. A)2 B)3 C)All real numbers D)No solutions

OpenStudy (anonymous):

First you need to simplify! distribute the -2's and continue from there.

OpenStudy (godlovesme):

Multiply -2 by x and 3 Then multiply -2 by x and 1

OpenStudy (godlovesme):

@infiniterosha u there?

OpenStudy (fibonaccichick666):

just like before, show us your work :)

OpenStudy (fibonaccichick666):

wait, it's the same problem...?

OpenStudy (anonymous):

Yeah its kinda the same thing. And sorry I was doing something

OpenStudy (anonymous):

-2(3 + x) = -2(x + 1) + -4 (3 * -2 + x * -2) = -2(x + 1) + -4 (-6 + -2x) = -2(x + 1) + -4

OpenStudy (anonymous):

-6 + -2x = -2(1 + x) + -4 -6 + -2x = (1 * -2 + x * -2) + -4 -6 + -2x = (-2 + -2x) + -4

OpenStudy (fibonaccichick666):

looks good to tyhere

OpenStudy (anonymous):

0=0

OpenStudy (fibonaccichick666):

how did you get that?

OpenStudy (fibonaccichick666):

I'm not saying it is wrong, but show me your work

OpenStudy (fibonaccichick666):

because if you stop a little earlier, the answer becomes clear

OpenStudy (anonymous):

Ok, because -6 + -2x = -2 + -4 + -2x -6 + -2x = -6 + -2x -6 + 6 + -2x = -6 + 6 + -2x

OpenStudy (fibonaccichick666):

ok, so stop at your second step there

OpenStudy (anonymous):

-6 + 6 = 0 0 + -2x = -6 + 6 + -2x -2x = -6 + 6 + -2x

OpenStudy (fibonaccichick666):

so go back to the second step here \[-6 + -2x = -6 + -2x\]

OpenStudy (fibonaccichick666):

pick an x, any x, will the equality hold?

OpenStudy (anonymous):

simplifying equations

OpenStudy (fibonaccichick666):

... uhm, random words, why are you saying that?

OpenStudy (anonymous):

okay there were gonna combine like terms.

OpenStudy (fibonaccichick666):

wait wait wait

OpenStudy (fibonaccichick666):

I mean literally stop here \[-6 + -2x = -6 + -2x\]

OpenStudy (anonymous):

okay.

OpenStudy (fibonaccichick666):

ok, now, pick an x any x, will the equality hold? ie will this always be true for any value of x

OpenStudy (anonymous):

3

OpenStudy (fibonaccichick666):

that is just a number... I need words. I have no clue what you are saying if you do not explain yourself, if you are saying "let's try x=3," then go ahead, see if it works

OpenStudy (anonymous):

I'm sorry someone was on my computer

OpenStudy (fibonaccichick666):

ohk